# An Explicit M1 Radiation-hydrodynamics Scheme for Three-dimensional Protostellar Evolution

### v1

@@ 2.2 Streaming/trapped component for an optically thick medium
which decreases as the optical depth increases. By contrast, the magnitude of the numerical diffusion is independent of the optical depth. For example, the Global Lax-Friedrichs (GLF) Riemann solver, which is used in this study for its simplicity and its ability to preserve symmetry more robustly than higher order schemes (J. Rosdahl et al., 2013), yields a numerical diffusion of the form
@@
To address this issue, J. Rosdahl & R. Teyssier (2015) proposed a method that reproduces the correct physical diffusion in the optically thick limit. Since the magnitude of the numerical diffusion scales with $E_{\mathrm{rad}}$, as shown in Equation (15), one can adjust $\bm{F}_{\mathrm{num}}$ to match the physical diffusion $\bm{F}_{\mathrm{diff}}$ by reducing $E_{\mathrm{rad}}$ in the RT calculation. In this approach, they split $E_{\mathrm{rad}}$ into two components, streaming photons $E_{\mathrm{rad,S}}$ and trapped photons $E_{\mathrm{rad,T}}$, and then use only $E_{\mathrm{rad,S}}$ for solving the advection terms. For the GLF solver, these components should be defined as
@@
Thus, we must modify the terms involving the radiation flux $\bm{F}$ in the RHD equations to include the flux contribution of $E_{\mathrm{rad,T}}$ as follows. The following treatment is essentially equivalent to that described in Appendix B of J. Rosdahl & R. Teyssier (2015), but without the impact of the RSLA, which we explicitly introduce later in Section 2.3. In the optically thick limit, the radiative flux in the comoving frame follows the diffusion approximation given by Equation (13). Meanwhile, the total flux in the lab frame includes advection by the moving gas that cannot be neglected in the dynamic diffusion regime as shown in Equation (A6). Accordingly, when $\tau_{\mathrm{cell}}\gg 1$ and the trapped component $E_{\mathrm{rad,T}}$ dominates, the radiative flux should be
@@
These equations are equivalent to those presented in Appendix B of J. Rosdahl & R. Teyssier (2015), except for some differences. First, we have not yet applied the RSLA to these equations. Next, we include scattering opacity, use lab frame quantities for the energy densities appearing in the radiation four-force terms, and retain only the dominant terms as described in Appendix B. Furthermore, we do not evolve the streaming and trapped energy densities separately, but instead describe the time evolution of their sum using Equation (26). This choice of evolving the combined quantity, rather than treating the two components individually, is merely a matter of formulation and does not represent a fundamental difference.
2.3 Reduced speed of light approximation
@@
Practically, we implement the RSLA by multiplying the advection terms by the reduction factor $f_{c}$, thereby reducing their magnitude. However, the specific formulation of the RSLA can vary depending on how $f_{c}$ is applied within the RHD equations (P. F. Hopkins et al., 2022). If the reduction is applied only to the advection terms, $\nabla\cdot\bm{F}$ and $c^{2}\nabla\cdot\bm{\mathsf{P}}_{\mathrm{rad,s}}$, which are responsible for the restrictive CFL condition, the resulting steady-state solutions may differ from those of the original, non-reduced system. To prevent this inconsistency, many previous studies have applied the reduction factor $f_{c}$ not only to the advection terms but also to the radiation source terms, $G^{0}$ and $\bm{G}$ (e.g., M. A. Skinner & E. C. Ostriker, 2013; J. Rosdahl & R. Teyssier, 2015; R. Kannan et al., 2019; P. Chang et al., 2020; H. Fukushima & H. Yajima, 2021). In this study, we adopt the same approach, except that we do not apply $f_{c}$ to the radiation advection term associated with gas dragging, $(4\bm{v}/3)E_{\mathrm{rad,T}}$ on the left-hand side of Equation (26). This is because this term represents the advection together with the gas and therefore does not impose a restrictive time step. Moreover, applying $f_{c}$ to this term would instead introduce an unphysical lag between radiation and gas motion, leading to unphysical behavior. With this treatment, Equations (26) and (27) are modified as follows:
@@
where $E^{\mathrm{R}}_{\mathrm{rad}}$ and $\bm{\mathsf{P}}^{\mathrm{R}}_{\mathrm{rad}}$ are identical to $E$ and $\bm{P}$ given by Equation (2) and (4) in J. Rosdahl & R. Teyssier (2015). With these variables, Equations (24), (25), (28)–(30), (33) and (34) can be rewritten as
@@
From Equations (36) and (38), it is evident that the sum $E_{\mathrm{gas}}+E^{\mathrm{R}}_{\mathrm{rad}}$ is conserved during the gas-radiation interaction under the RSLA. The radiation pressure tensor $\bm{\mathsf{P}}^{\mathrm{R}}_{\mathrm{rad,S}}$ is calculated from $E^{\mathrm{R}}_{\mathrm{rad,S}}$ and $F_{\mathrm{S}}$ as
@@
3 Challenges of RHD simulations that resolve the protostellar interior
In this section, we discuss the three key challenges encountered when applying the M1 closure method with the RSLA to star formation RHD simulations that resolve the protostellar interior. We present solutions to each of these challenges in Section 4.
3.1 Challenge 1: energy conservation with the RSLA
While the formulation of the RSLA introduced in Section 2.3 is widely adopted to improve computational efficiency, it induces the non-conservation of total physical energy. Actually, from Equations (33) and (36), the change of the sum of the gas and radiation energy $E_{\mathrm{gas}}+E_{\mathrm{rad}}$ during gas–radiation interactions can be described by
$\frac{\partial}{\partial t}(E_{\mathrm{gas}}+E_{\mathrm{rad}})=cG^{0}(1-f_{c}),$ (44)
which implies that physical energy is effectively lost from the system whenever it is transferred from the gas to the radiation with $G^{0}<0$.
@@ 3.1 Challenge 1: energy conservation with the RSLA
In contrast, in optically thick systems, when the temperature is sufficiently high, the radiation energy becomes dominant within the medium. When this energy is artificially lost due to the RSLA, we cannot accurately follow the thermal and dynamical evolution of the system. In Section 4.1, we introduce a new scheme designed to restore physical energy conservation in the optically thick limit.
@@
3.2 Challenge 2: radiative transfer with steep optical depth gradient
@@
As mentioned in Section 2.2, the original M1 closure scheme fails to solve RT in regions where the optical depth per cell exceeds unity due to excessive numerical diffusion. Then, to accurately follow the RT in such a regime, we introduce the method by J. Rosdahl & R. Teyssier (2015), which splits the radiation energy $E^{\mathrm{R}}_{\mathrm{rad}}$ into streaming component $E^{\mathrm{R}}_{\mathrm{rad,S}}$ and trapped component $E^{\mathrm{R}}_{\mathrm{rad,T}}$. However, even with this method, a problem in RT still remains.
@@
As an example, consider a situation in which an optically thick cell with $\tau_{\mathrm{cell}}\gg 1$ is directly adjacent to an optically thin cell with $\tau_{\mathrm{cell}}\ll 1$. Physically, radiation energy should escape from the surfaces of the optically thick cell at a flux of $\sigma_{\mathrm{SB}}T^{4}$ per unit area per unit time. However, when we use the method introduced in Section 2.2, the trapped component $E^{\mathrm{R}}_{\mathrm{rad,T}}$ dominates and the streaming component $E^{\mathrm{R}}_{\mathrm{rad,S}}$ is suppressed to very small values in optically thick cells. Since we calculate RT only with $E^{\mathrm{R}}_{\mathrm{rad,S}}$, radiation cannot effectively escape from the optically thick cell.
@@
This situation is particularly relevant at the surface of a protostar, where the density and opacity vary steeply over short spatial scales. For a solar-mass main-sequence star, this scale is only on the order of $10^{2}$ km. These scales are far too small to be resolved in simulations that follow the global evolution of the entire star. However, accurately solving RT across such boundaries is essential, because it determines how efficiently the protostar loses entropy through radiation and how much radiative feedback is imparted to the surrounding gas. We address this issue in Section 4.2.
3.3 Challenge 3: modeling the radiation spectrum
@@
Since the gas opacity to radiation is highly sensitive to frequency, the interaction between gas and radiation, such as photoheating, photoionization, and radiation pressure, depends strongly on the shape of the radiation spectrum. In this study, we also estimate the mean opacity with the spectrum as in Equations (9) and (10). While the most straightforward method for accurately following the evolution of the spectrum is solving the RHD equations for many frequency bins, their computational cost scales linearly with the number of bins and becomes prohibitively expensive. Instead, most RHD simulations employ only one or a few frequency bins, and assume a spectral shape within each bin.
@@
Alternatively, in simulations that employ sink particles to represent individual stars or stellar clusters, the spectral shape is often determined based on 1D stellar evolution models. However, such an approach is not applicable to simulations like ours, in which stars are resolved directly during the course of the calculation and the radiation spectrum must be determined consistently with their 3D evolution. Moreover, the above method cannot capture spatial variations in the radiation spectrum. This challenge is not specific to the M1 scheme but arises with any closure relation, including the variable Eddington tensor method, and is therefore a general difficulty of RHD simulations. In Section 4.3, we describe our method for capturing the time evolution and spatial variation of the spectral shape.
4 Improved Scheme
@@
In this section, we present our solutions to the challenges outlined in Section 3. Section 4.1 describes how we address the violation of energy conservation caused by the RSLA. In Section 4.2, we introduce a method for accurately solving RT in systems with a steep optical depth gradient. Finally, in Section 4.3, we describe our approach for evaluating the radiation spectrum.
4.1 Hybrid radiative transfer: RSLA and Non-RSLA components

### v2

@@ II.2 Streaming/trapped component for an optically thick medium
which decreases as the optical depth increases. By contrast, the magnitude of the numerical diffusion is independent of the optical depth. For example, the Global Lax-Friedrichs (GLF) Riemann solver, which is used in this study for its simplicity and its ability to preserve symmetry more robustly than higher order schemes (Rosdahl et al., 2013), yields a numerical diffusion of the form
@@
To address this issue, Rosdahl and Teyssier (2015) proposed a method that reproduces the correct physical diffusion in the optically thick limit. Since the magnitude of the numerical diffusion scales with $E_{\mathrm{rad}}$, as shown in Equation (15), one can adjust $\bm{F}_{\mathrm{num}}$ to match the physical diffusion $\bm{F}_{\mathrm{diff}}$ by reducing $E_{\mathrm{rad}}$ in the RT calculation. In this approach, they split $E_{\mathrm{rad}}$ into two components, streaming photons $E_{\mathrm{rad,S}}$ and trapped photons $E_{\mathrm{rad,T}}$, and then use only $E_{\mathrm{rad,S}}$ for solving the advection terms. For the GLF solver, these components should be defined as
@@
Thus, we must modify the terms involving the radiation flux $\bm{F}$ in the RHD equations to include the flux contribution of $E_{\mathrm{rad,T}}$ as follows. The following treatment is essentially equivalent to that described in Appendix B of Rosdahl and Teyssier (2015), but without the impact of the RSLA, which we explicitly introduce later in Section II.3. In the optically thick limit, the radiative flux in the comoving frame follows the diffusion approximation given by Equation (13). Meanwhile, the total flux in the lab frame includes advection by the moving gas that cannot be neglected in the dynamic diffusion regime as shown in Equation (A6). Accordingly, when $\tau_{\mathrm{cell}}\gg 1$ and the trapped component $E_{\mathrm{rad,T}}$ dominates, the radiative flux should be
@@
These equations are equivalent to those presented in Appendix B of Rosdahl and Teyssier (2015), except for some differences. First, we have not yet applied the RSLA to these equations. Next, we include scattering opacity, use lab frame quantities for the energy densities appearing in the radiation four-force terms, and retain only the dominant terms as described in Appendix B. Furthermore, we do not evolve the streaming and trapped energy densities separately, but instead describe the time evolution of their sum using Equation (26). This choice of evolving the combined quantity, rather than treating the two components individually, is merely a matter of formulation and does not represent a fundamental difference.
II.3 Reduced speed of light approximation
@@
Practically, we implement the RSLA by multiplying the advection terms by the reduction factor $f_{c}$, thereby reducing their magnitude. However, the specific formulation of the RSLA can vary depending on how $f_{c}$ is applied within the RHD equations (Hopkins et al., 2022). If the reduction is applied only to the advection terms, $\nabla\cdot\bm{F}$ and $c^{2}\nabla\cdot\bm{\mathsf{P}}_{\mathrm{rad,S}}$, which are responsible for the restrictive CFL condition, the resulting steady-state solutions may differ from those of the original, non-reduced system. To prevent this inconsistency, many previous studies have applied the reduction factor $f_{c}$ not only to the advection terms but also to the radiation source terms, $G^{0}$ and $\bm{G}$ (e.g., Skinner and Ostriker, 2013; Rosdahl and Teyssier, 2015; Kannan et al., 2019; Chang et al., 2020; Fukushima and Yajima, 2021). In this study, we adopt the same approach, except that we do not apply $f_{c}$ to the radiation advection term associated with gas dragging, $(4\bm{v}/3)E_{\mathrm{rad,T}}$ on the left-hand side of Equation (26). This is because this term represents the advection together with the gas and therefore does not impose a restrictive time step. Moreover, applying $f_{c}$ to this term would instead introduce an unphysical lag between radiation and gas motion, leading to unphysical behavior. With this treatment, Equations (26) and (27) are modified as follows:
@@
where, from Equation (16), $E^{\mathrm{R}}_{\mathrm{rad}}$, $E^{\mathrm{R}}_{\mathrm{rad,S}}$, and $E^{\mathrm{R}}_{\mathrm{rad,T}}$ are related as
$\begin{array}[]{l}E^{\mathrm{R}}_{\mathrm{rad,S}}=(1-\eta_{\mathrm{T}})E^{\mathrm{R}}_{\mathrm{rad}},\hskip 8.53581ptE^{\mathrm{R}}_{\mathrm{rad,T}}=\eta_{\mathrm{T}}E^{\mathrm{R}}_{\mathrm{rad}},\\[6.0pt] {\displaystyle\eta_{\mathrm{T}}=\exp\left(-\frac{2}{3\tau_{\mathrm{cell}}}\right).}\end{array}$ (36)
Here, $E^{\mathrm{R}}_{\mathrm{rad}}$ and $\bm{\mathsf{P}}^{\mathrm{R}}_{\mathrm{rad}}$ are identical to $E$ and $\bm{P}$ given by Equation (2) and (4) in Rosdahl and Teyssier (2015). With these variables, Equations (24), (25), (28)–(30), (33) and (34) can be rewritten as
@@
From Equations (37) and (39), it is evident that the sum $E_{\mathrm{gas}}+E^{\mathrm{R}}_{\mathrm{rad}}$ is conserved during the gas-radiation interaction under the RSLA. The radiation pressure tensor $\bm{\mathsf{P}}^{\mathrm{R}}_{\mathrm{rad,S}}$ is calculated from $E^{\mathrm{R}}_{\mathrm{rad,S}}$ and $\bm{F}_{\mathrm{S}}$ as
@@
III Challenges of RHD simulations that resolve the protostellar interior
In this section, we discuss the three key challenges encountered when applying the M1 closure method with the RSLA to star formation RHD simulations that resolve the protostellar interior. We present solutions to each of these challenges in Section IV.
III.1 Challenge 1: energy conservation with the RSLA
While the formulation of the RSLA introduced in Section II.3 is widely adopted to improve computational efficiency, it induces the non-conservation of total physical energy. Actually, as can be seen from Equations (33) and (37), when $G^{0}<0$, the gas loses an amount of energy $cG^{0}$ through its interaction with radiation, whereas the radiation gains only $f_{c}cG^{0}$. This implies that, whenever the energy is transferred from the gas to the radiation with $G^{0}<0$, the radiation field receives only a fraction $f_{c}$ of the energy it would physically acquire, and the remaining fraction $(1-f_{c})$ is artificially removed from the total energy budget.
@@ III.1 Challenge 1: energy conservation with the RSLA
In contrast, in optically thick systems, when the temperature is sufficiently high, the radiation energy becomes dominant within the medium. When this energy is artificially lost due to the RSLA, we cannot accurately follow the thermal and dynamical evolution of the system (see, e.g., the test problem in Wibking and Krumholz, 2022). In Section IV.1, we introduce a new scheme designed to restore physical energy conservation in the optically thick limit.
@@
III.2 Challenge 2: radiative transfer with steep optical depth gradient
@@
As mentioned in Section II.2, the original M1 closure scheme fails to solve RT in regions where the optical depth per cell exceeds unity due to excessive numerical diffusion. Then, to accurately follow the RT in such a regime, we introduce the method by Rosdahl and Teyssier (2015), which splits the radiation energy $E^{\mathrm{R}}_{\mathrm{rad}}$ into streaming component $E^{\mathrm{R}}_{\mathrm{rad,S}}$ and trapped component $E^{\mathrm{R}}_{\mathrm{rad,T}}$. However, even with this method, a problem in RT still remains.
@@
As an example, consider a situation in which an optically thick cell with $\tau_{\mathrm{cell}}\gg 1$ is directly adjacent to an optically thin cell with $\tau_{\mathrm{cell}}\ll 1$. Physically, radiation energy should escape from the surfaces of the optically thick cell at a flux of $\sigma_{\mathrm{SB}}T^{4}$, where $\sigma_{\mathrm{SB}}$ is the Stefan–Boltzmann constant. However, when we use the method introduced in Section II.2, the trapped component $E^{\mathrm{R}}_{\mathrm{rad,T}}$ dominates and the streaming component $E^{\mathrm{R}}_{\mathrm{rad,S}}$ is suppressed to very small values in optically thick cells. Since we calculate RT only with $E^{\mathrm{R}}_{\mathrm{rad,S}}$, radiation cannot effectively escape from the optically thick cell.
@@
This situation is particularly relevant at the surface of a protostar, where the density and opacity vary steeply over short spatial scales. For a solar-mass main-sequence star, this scale is only on the order of $10^{2}$ km. These scales are far too small to be resolved in simulations that follow the global evolution of the entire star. However, accurately solving RT across such boundaries is essential, because it determines how efficiently the protostar loses entropy through radiation and how much radiative feedback is imparted to the surrounding gas. We address this issue in Section IV.2.
III.3 Challenge 3: modeling the radiation spectrum
@@
Since the gas opacity to radiation is highly sensitive to frequency, the interaction between gas and radiation, such as photoheating, photoionization, and radiation pressure, depends strongly on the shape of the radiation spectrum. In this study, we also estimate the mean opacity with the spectrum as in Equations (9) and (10). While the most straightforward method for accurately following the evolution of the spectrum is solving the RHD equations for many frequency bins, their computational cost scales linearly with the number of bins and becomes prohibitively expensive. Instead, most RHD simulations employ only one (grey approximation) or a few frequency bins, and assume a spectral shape within each bin (Skinner and Ostriker, 2013; Mignon-Risse et al., 2020; Wibking and Krumholz, 2022).
@@
Alternatively, in simulations that employ sink particles to represent individual stars or stellar clusters, the spectral shape is often determined based on 1D stellar evolution models. However, such an approach is not applicable to simulations like ours, in which stars are resolved directly during the course of the calculation and the radiation spectrum must be determined consistently with their 3D evolution. Moreover, the above method cannot capture spatial variations in the radiation spectrum. This challenge is not specific to the M1 scheme but arises with any closure relation, including the variable Eddington tensor method, and is therefore a general difficulty of RHD simulations. In Section IV.3, we describe our method for capturing the time evolution and spatial variation of the spectral shape.
IV Improved Scheme
@@
In this section, we present our solutions to the challenges outlined in Section III. Section IV.1 describes how we address the violation of energy conservation caused by the RSLA. In Section IV.2, we introduce a method for accurately solving RT in systems with a steep optical depth gradient. Finally, in Section IV.3, we describe our approach for evaluating the radiation spectrum. For the convenience of the readers, before proceeding to the description of the improved scheme, we first summarize the notation and relations for the constants, gas variables, and radiation variables used in the improved scheme in Tables 1, 2, and 3.
Table 1: Notation for constants
Symbol Meaning
$k_{\mathrm{B}}$ Boltzmann constant
$m_{\mathrm{H}}$ mass of a hydrogen nucleus
$h$ Planck constant
$\sigma_{\mathrm{SB}}$ Stefan–Boltzmann constant
$a$ radiation constant
$G$ gravitational constant
$c$ speed of light
$f_{c}$ reduction factor in the RSLA
$\tilde{c}=f_{c}c$ reduced speed of light
$\Delta x$ computational cell width
$\Delta t$ computational time step
Table 2: Notation for gas variables
Symbol Meaning
$\rho$ gas density
$P_{\mathrm{gas}}$ gas pressure
T gas temperature
$\bm{v}$ gas velocity
$\beta=|\bm{v}|/c$ ratio of gas speed to light speed
$c_{\mathrm{s}}$ sound speed
$\mu$ mean molecular weight
$\gamma$ specific heat ratio
$E_{\mathrm{gas}}$ gas total energy
$\phi$ gravitational potential
$\bm{g}$ gravitational acceleration
Table 3: Notation and relations for radiation variables. For simplicity, variables related to the photon number density are omitted.
Symbol Relation Meaning
Energy-density variables
$E^{\mathrm{con}}_{\mathrm{rad}}$ $E^{\mathrm{con}}_{\mathrm{rad}}=E^{\mathrm{R}}_{\mathrm{rad}}+E^{\mathrm{N}}_{\mathrm{rad}}$ total conserved radiation energy density variable
$E^{\mathrm{R}}_{\mathrm{rad}}$ $E^{\mathrm{R}}_{\mathrm{rad}}=(1-\eta^{\mathrm{N}})E^{\mathrm{con}}_{\mathrm{rad}}=E^{\mathrm{R,S}}_{\mathrm{rad}}+E^{\mathrm{R,T}}_{\mathrm{rad}}$ reduced speed of light approximation (RSLA) component
(scaled by $1/f_{c}$ from the physical energy density)
$E^{\mathrm{N}}_{\mathrm{rad}}$ $E^{\mathrm{N}}_{\mathrm{rad}}=\eta^{\mathrm{N}}E^{\mathrm{con}}_{\mathrm{rad}}$ non-RSLA (full speed of light) component
(corresponding to the physical energy density)
$E^{\mathrm{R,S}}_{\mathrm{rad}}$ $E^{\mathrm{R,S}}_{\mathrm{rad}}=(1-\eta_{\mathrm{T}})E^{\mathrm{R}}_{\mathrm{rad}}$ streaming component of the RSLA radiation field
$E^{\mathrm{R,T}}_{\mathrm{rad}}$ $E^{\mathrm{R,T}}_{\mathrm{rad}}=\eta_{\mathrm{T}}E^{\mathrm{R}}_{\mathrm{rad}}$ trapped component of the RSLA radiation field
Fraction and optical-depth per computational cell
$\eta^{\mathrm{N}}$ $\eta^{\mathrm{N}}=\exp\!\left(-\frac{2}{3f_{c}\tau_{\mathrm{cell}}}\right)$ non-RSLA fraction of total radiation field
$\eta_{\mathrm{T}}$ $\eta_{\mathrm{T}}=\exp\!\left(-\frac{2}{3\tau_{\mathrm{cell}}}\right)$ trapped fraction of the RSLA radiation field
$\tau_{\mathrm{cell}}$ $\tau_{\mathrm{cell}}=\rho\chi_{\mathrm{0R}}\Delta x$ optical depth per computational cell
RSLA-streaming transport and closure variables
$\bm{F}_{\mathrm{S}}$ radiation flux of the RSLA-streaming component
$\bm{\mathsf{P}}^{\mathrm{R}}_{\mathrm{rad,S}}$ $\bm{\mathsf{P}}^{\mathrm{R}}_{\mathrm{rad,S}}=\bm{\mathsf{D}}_{\mathrm{S}}E^{\mathrm{R}}_{\mathrm{rad,S}}$ radiation pressure tensor of the RSLA-streaming component
$\bm{\mathsf{D}}_{\mathrm{S}}$ Eddington tensor for the RSLA-streaming component
Radiation four-force density and decomposition
$G^{0}$ time component of radiation four-force density
$\bm{G}_{\mathrm{tot}}$ $\bm{G}_{\mathrm{tot}}=\bm{G}^{\mathrm{R}}_{\mathrm{S}}+\bm{G}^{\mathrm{R}}_{\mathrm{T}}+\bm{G}^{\mathrm{N}}$ total spatial component of radiation four-force density
$\bm{G}^{\mathrm{R}}_{\mathrm{S}}$ RSLA-streaming component of $\bm{G}_{\mathrm{tot}}$
$\bm{G}^{\mathrm{R}}_{\mathrm{T}}$ RSLA-trapped component of $\bm{G}_{\mathrm{tot}}$
$\bm{G}^{\mathrm{N}}$ non-RSLA component of $\bm{G}_{\mathrm{tot}}$
Opacity and spectral variables
$\kappa^{\nu}_{0}$ frequency-dependent absorption opacity in the comoving frame
$\sigma^{\nu}_{0}$ frequency-dependent scattering opacity in the comoving frame
$\kappa_{\mathrm{0P}}$ Planck mean absorption opacity
$\kappa_{0E}$ energy-mean absorption opacity
$\chi_{0E}$ energy-mean total (absorption + scattering) opacity
$\chi_{0F}$ flux-mean total opacity
$\chi_{\mathrm{0R}}$ Rosseland mean opacity
$\chi_{\mathrm{crit}}$ critical opacity above which the non-RSLA component dominates
$\hat{\chi}$ $\hat{\chi}=\max(\chi_{0R},\chi_{\mathrm{crit}})$ effective opacity for diffusive transport of the non-RSLA component
$B^{\nu}(T)$ Planck function at temperature $T$
$E^{\nu}_{\mathrm{rad},0}$ comoving-frame radiation energy density per unit frequency
$F^{\nu}_{0}$ comoving-frame radiation flux per unit frequency
$T_{\mathrm{rad}}$ radiation temperature defined from the mean photon energy
IV.1 Hybrid radiative transfer: RSLA and Non-RSLA components